OpenStudy (anonymous):

PLEASE HELP!!!! There are nickels and dimes in a candy jar. If there are 22 more nickels than dimes write an expression that represents of the total number of coins if the number of dimes is d.

OpenStudy (anonymous):

well how much do nickels cost

OpenStudy (anonymous):

in value -.-

OpenStudy (anonymous):

and dimes also

OpenStudy (kainui):

It almost appears that you forgot how much nickels and dimes are worth @iforgot lol

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

lol well there was no reply so apparently it was not me

OpenStudy (anonymous):

nickels are 5 and dimes are 10 ?

OpenStudy (anonymous):

cool give them each a variable but since you said

OpenStudy (anonymous):

scratch "since u said"

OpenStudy (anonymous):

5n

OpenStudy (anonymous):

10d

OpenStudy (anonymous):

there is a jar you have 22 nickel MORE THAN dimes

OpenStudy (anonymous):

total number coins = 22nickels MORE THAN dimes o forgot you also need a variable for total amount or coins in jar how would you put that in an equation

OpenStudy (anonymous):

choose your variable for total amount of coins in jar MORE THAN (in math ) denotes what?

OpenStudy (anonymous):

addding

OpenStudy (anonymous):

It's 4 times than 2 less than the answer to 4+5 time 3 to the third power which divides into 12.

OpenStudy (anonymous):

so whats the answer ?

OpenStudy (anonymous):

yep ill chose a variable for you than :P total amount of coins = A total number coins = 22nickels plus dimes

OpenStudy (anonymous):

are you able to write out the expression

OpenStudy (anonymous):

you kinda found out the answer just go over it

OpenStudy (anonymous):

22n+10d ?

OpenStudy (anonymous):

lol what the heck i think i did the whole thing wrong in the first place

OpenStudy (anonymous):

it only asks for the number of coins

OpenStudy (anonymous):

very stupid mistake XD

OpenStudy (anonymous):

lol omg

OpenStudy (anonymous):

how would you do that then ?

OpenStudy (anonymous):

the coins just dont have values

OpenStudy (anonymous):

i meant the logic was big mistake

OpenStudy (anonymous):

still done the same way

OpenStudy (anonymous):

so u wouldnt write the 5 and 10

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

soo 22>d ? idk

OpenStudy (texaschic101):

n = d + 22 .10d + .05n = $ amount of coins in jar sun d + 22 in for n .10d + .05(d + 22) = total number of coins (T) .10d + .05d + 1.1 = T .15d = T - 1.1 d = (T - 1.1)/.15 you probably need a second opinion on this :/

OpenStudy (texaschic101):

oops typo....sub not sun..lol

OpenStudy (anonymous):

u only need an expression :P

OpenStudy (texaschic101):

then I guess the expression is : .10d + .05(d + 22) = T

OpenStudy (anonymous):

an expression Does NOT need = marks

OpenStudy (texaschic101):

then leave the = T off.......idk

OpenStudy (anonymous):

how sure are u this is right ?

OpenStudy (anonymous):

texas is more correct ^^

OpenStudy (texaschic101):

Lets just say, that is what I would put......you can get a 2nd opinion if you want to

OpenStudy (anonymous):

huh

OpenStudy (anonymous):

im a go with ur answer texas

OpenStudy (anonymous):

start out saying u have at least 22 and being more than 22 a

OpenStudy (anonymous):

also*

OpenStudy (anonymous):

Johan's age is given by the polynomial x2 - 3x + 7. Bret is younger and his age is given by the polynomial x2 - x + 1. What is difference in their ages?

OpenStudy (texaschic101):

there is 22 more nickels then dimes ...so n = d + 22 dimes are .10d nickels are .05n .10d + .05(d + 22) -- I subbed d + 22 in for n do you understand what I did ?

OpenStudy (anonymous):

one thing about texas' answer its only counting # of coins not value

OpenStudy (texaschic101):

thats true

OpenStudy (anonymous):

It looks right

OpenStudy (texaschic101):

x^2 - 3x + 7 x^2 - x + 1 ------------subtract -4x + 8 ??? I am not sure about this one

OpenStudy (anonymous):

yes plz

OpenStudy (anonymous):

u come bearing gifts to be received gratefully at least by me :D

OpenStudy (anonymous):

flattery ^^

OpenStudy (texaschic101):

can you do the question with polynomials because I don't know if I did that one correctly